# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v3

@@ Abstract
Text-driven 3D indoor scene generation is useful for gaming, the film industry, and AR/VR applications. However, existing methods cannot faithfully capture the room layout, nor do they allow flexible editing of individual objects in the room. To address these problems, we present Ctrl-Room, which can generate convincing 3D rooms with designer-style layouts and high-fidelity textures from just a text prompt. Moreover, Ctrl-Room enables versatile interactive editing operations such as resizing or moving individual furniture items. Our key insight is to separate the modeling of layouts and appearance. Our proposed method consists of two stages: a Layout Generation Stage and an Appearance Generation Stage. The Layout Generation Stage trains a text-conditional diffusion model to learn the layout distribution with our holistic scene code parameterization. Next, the Appearance Generation Stage employs a fine-tuned ControlNet to produce a vivid panoramic image of the room guided by the 3D scene layout and text prompt. We thus achieve a high-quality 3D room generation with convincing layouts and lively textures. Benefiting from the scene code parameterization, we can easily edit the generated room model through our mask-guided editing module, without expensive edit-specific training. Extensive experiments on the Structured3D dataset demonstrate that our method outperforms existing methods in producing more reasonable, view-consistent, and editable 3D rooms from natural language prompts.
Keywords:
3D Indoor Scene Generation Layout generation panorama generation
@@ 1 Introduction
High-quality textured 3D models play a crucial role across a wide array of applications, ranging from interior design and video games to simulators for embodied AI. Among all types of 3D content, indoor scenes hold particular interest due to their complexity and utility. Traditionally, 3D indoor scenes are crafted manually by professional artists, which is both time-consuming and costly. While recent advancements in generative models [23, 5, 15, 30] have simplified the creation of 3D models from textual descriptions, extending this capability to text-driven 3D indoor scene generation presents unique challenges as they exhibit strong semantic layout constraints, such as neighboring walls are perpendicular and the TV set often faces a sofa, that are more complicated than objects.
@@
Existing text-driven 3D indoor scene generation approaches, such as Text2-Room [12] and Text2NeRF [47], are designed with an incremental framework. They create 3D indoor scenes by incrementally generating different viewpoints frame-by-frame and reconstructing the 3D mesh of the room from these sub-view images. However, these approaches often fail to model the global layout of the room, resulting in unconvincing results that lack semantic plausibility. As shown in the left of  Fig. 1 (a), the result of Tex2Room exhibits repeating objects, e.g. several cabinets in a living room, and does not follow the furniture layout patterns. We refer to this problem as the ‘Penrose Triangle problem’, where a generated scene has plausible 3D structures everywhere locally but lacks global consistency. Furthermore, prior approaches do not offer user-friendly interactive manipulation, as the resulting 3D geometry and textures are not editable.
Indoor scenes might also be represented by a panorama image. Several works [17, 18, 33, 39] have been proposed to generate such a panorama from a text prompt. We might further recover the depth map of these images to build a textured 3D room model. However, these works cannot guarantee correct room layouts. As shown on the right side of Fig. 1 (a), a bedroom generated by MVDiffusion [39] contains multiple beds, which violates room layout priors. Furthermore, these methods cannot easily control the individual objects.
@@
To address these shortcomings, we propose a novel two-stage method to generate a high-fidelity and editable 3D room. The key insight is to separate the generation of 3D geometric layouts from that of visual appearance, which allows us to better capture the room layout and achieve flexible editing at the same time. In the first stage, from text input, our method creates plausible room layouts with various furniture types and positions. Unlike previous scene synthesis methods  [37, 22] that only focus on the furniture arrangement, our approach further considers walls with doors and windows, which play an essential role in the layout. To achieve this goal, we parameterize the room by a holistic scene code, which represents a room as a set of objects. Each object is represented by a vector capturing its position, size, semantic class, and orientation. Based on our compact parameterization, we design a diffusion model to learn the 3D room layout distribution from the Structured3D dataset [49].
@@
Our method then generates the room appearance with the guidance of the 3D room layout. We model the appearance of the room as a panoramic image, which is generated by a text-to-image latent diffusion model. Unlike previous text-to-panorama works  [39, 6], our method explicitly enforces room layout constraints and guarantees plausible 3D room structures and furniture arrangement. To achieve this goal, we convert the 3D layout synthesized in the first stage into a semantic segmentation map and feed it to a fine-tuned ControlNet [48] model to create the panorama image.
@@
Most importantly, benefiting from the separation of layout and appearance, our method enables flexible editing on the generated 3D room. The user can replace or modify the size and position of furniture items, e.g. replacing the TV and TV stand or moving up the picture as in Fig. 1 (b), by instructions or mouse clicks. Our method can update the room according to the edited room layout through our mask-guided editing module without expensive edit-specific training. The updated room appearance maintains consistency with the original version while satisfying the user’s edits. To our knowledge, it’s the first work that achieves 3D indoor scene editing through a 2D diffusion model without expensive edit-specific training.
@@
  • To address the Penrose Triangle Problem, we design a two-stage method for 3D room generation from pure text input, which separates the geometric layout generation and appearance generation. In this way, our method can better capture the room layout constraints in real-world data and produce a vivid and rich appearance at the same time.
  • Our separation of geometric layout and visual appearance allows us to have flexible control and editing over the generated 3D room model. Users can adjust the size, semantic class, and position of furniture items easily.
  • We introduce a novel method to generate and edit panoramic images, which achieves high-quality results with loop consistency through a pre-trained latent image diffusion model without expensive edit-specific training.
@@ 2.1 Text-based 3D Object Generation
Early methods employ 3D datasets to train generative models. Text2Shape [4] learns a feature representation from paired text and 3D data and uses GAN to generate 3D shapes from text. Point-E [20] and Shap-E [13] enlarge the the scope of the training dataset and employ a latent diffusion model [25] for object generation. However, 3D datasets are scarce which makes these methods difficult to scale. More recent methods [20, 23, 15, 41, 5, 43] exploit the powerful 2D text-to-image diffusion models [25, 27] for 3D model generation. Typically, these methods generate one or multiple 2D images in an incremental fashion and optimize the 3D model accordingly. DreamFusion [23] introduces a loss based on probability density distillation and optimizes a randomly initialized 3D model through gradient descent. Magic3D [15] uses a coarse model to represent 3D content and accelerates it using a sparse 3D hash grid structure. Score Jacobian Chaining [41] aggregates the results of the 2D diffusion models to generate a 3D scene. Fantasia3D [5] optimizes a mesh from scratch with DMTet [31] and stable diffusion [25]. To alleviate over-saturation, over-smoothing, and low-diversity problems, ProlificDreamer [43] models and optimizes the 3D parameters, NeRF [19] or mesh, through variational score distillation. However, all these methods focus on text-based 3D object generation. They cannot be directly applied to create 3D rooms that have additional structural layout constraints.
@@ 2.2 Text-based 3D Room Generation
Layout generation has been greatly boosted by trans-former-based methods. LayoutTransformer [10] employs self-attention to capture relationships between elements to accomplish layout completion. ATISS [22] proposes an autoregressive transformer to generate proper indoor scenes with only the room type and floor plan as the input. DiffuScene [37] and Instruct-Scene [16] model a union of furniture as a fully connected scene graph and propose a diffusion model to sample physically plausible scenes. While these methods generate reasonable furniture layouts, they do not consider the walls, doors, and windows which are crucial in the arrangement of the furniture. Therefore, these methods do not always generate realistic indoor environments.
@@
Another line of work [17, 18, 33] represent an indoor scene by a panorama image without modeling 3D shapes. These methods enjoy the benefits of abundant image training data and produce vivid results. COCO-GAN [17] produces a set of patches and assemble them into a panoramic image. InfinityGAN [18] uses the information of two patches to generate the parts between them, to finally obtain a panoramic image.  [33] proposes a 360-aware layout generator to produce furniture arrangements and uses this layout to synthesize a panoramic image based on the input scene background. MVDiffusion [39] simultaneously generates multi-view perspective images and proposes a correspondence-aware attention block to maintain multi-view consistency, and then transfers these images to a panorama. These methods might suffer from incorrect room layout since they do not enforce layout constraints. Furthermore, the results of these methods cannot be easily edited, e.g. resizing or moving furniture around, because they do not maintain an object-level representation.
@@
GAUDI [2] generates immersive 3D indoor scenes rendered from a moving camera. It disentangles the 3D representation and camera poses to ensure the consistency of the scene during camera movement. CC3D [1] proposes a 3D-aware GAN for multi-object scenes conditioned on a single semantic layout image and is trained using posed multi-view RGB images. Another related line of work [35, 45, 29] deals with retexturizing a given representation of 3D scenes. They employ 2D diffusion models to stylize and further improve the given geometry. Text2Room [12] incrementally synthesizes nearby images with a 2D diffusion model and recovers its depth maps to stitch these images into a 3D room model. Text2Room is the closest to our work, but it cannot handle the geometric and textural consistency among the images, resulting in the ‘Penrose Triangle problem’. In our method, we take both geometry and appearance into consideration and create a more geometrically plausible 3D room.
@@ 3 Method
Figure 2: Overview of our method. In the Layout Generation Stage, we synthesize a scene code from the text input and convert it to a 3D bounding box representation to facilitate editing. In the Appearance Generation Stage, we project the bounding boxes into a semantic segmentation map to guide the panorama synthesis. The panorama is then reconstructed into a textured 3D mesh model.
In order to achieve text-based 3D indoor scene generation and editing, we propose Ctrl-Room. We first generate room layout from an input text and then generate the room appearance according to the layout, followed by panoramic reconstruction to generate the final 3D textured mesh. This mechanism solves the Penrose Triangle Problem to generate physically plausible 3D rooms, while also enabling users to edit the scene layout interactively. The overall framework of our method is depicted in  Fig. 2, which consists of two stages: the Layout Generation Stage and the Appearance Generation Stage. In the Layout Generation Stage, we parameterize the indoor scene with a holistic scene code and design a diffusion model to learn its distribution. Once the holistic scene code is generated from text, we recover the room as a set of orientated bounding boxes of walls and objects. Note that users can edit these bounding boxes by dragging objects to adjust their semantic types, positions, or scales, enabling the customization of 3D scene results according to the user’s preferences. In the Appearance Generation Stage, we obtain an RGB panorama through a conditioned image diffusion model to represent the room texture. Specifically, we project the generated layout bounding boxes into a semantic segmentation layout. We then fine-tune a pre-trained ControlNet [48] model to generate an RGB panorama from the input room layout. To ensure loop consistency, we propose a novel loop-consistent sampling during the inference process. Finally, the textured 3D mesh is obtained by estimating the depth map of the generated panorama.
@@ 3.1 Layout Generation Stage
Scene Code Definition. Different from previous methods  [22, 37], we consider not only furniture but also walls, doors, and windows to define the room layout. We employ a unified encoding of various objects. Specifically, given a 3D scene $\mathcal{S}$ with $m$ walls and $n$ furniture items, we represent the scene layout as a holistic scene code $\mathbf{x_{0}}=\{\mathbf{o_{i}}\}^{N}_{i=1}$, where $N=m+n$. We encode each object $o_{j}$ as a node with various attributes, viz., center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. The concatenation of these attributes characterizes each node as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},r_{i}]$. As can be seen in Fig. 3 (a), we represent a scene layout as a tensor $\mathbf{x_{0}}\in\mathbb{R}^{N\times D}$, where $D$ is the attribute dimension of a node. In all the data, we choose the normal direction of the largest wall as the ‘main direction’. For other objects, we take the angles between their front directions and the main direction as their rotations. We use the one-hot encoding to represent their semantic types, such as sofa or lamp. For more details on the scene encoding, please refer to the Appendix.
@@
$\mathcal{L}_{\rm\mathbf{w-o}}=\sum_{i=1}^{K_{\rm wall}}\sum_{j=1}^{K_{\rm object}}\sum_{p=1}^{8}Relu[-(a_{i}x_{jp}+b_{i}y_{jp}+c_{i}z_{jp}+d_{i})]\mathds{1}(\prod_{\mathbf{w}_{i}}(x_{jp},y_{jp},z_{jp})\,{\rm in}\,\mathbf{w}_{i}).$ (5)
Here, $(a_{i},b_{i},c_{i})$ is the normal vector of wall $\mathbf{w}_{i}$ that points towards the room center. $\prod_{\mathbf{w}_{i}}$ is the operator projecting a point onto the plane defined by $\mathbf{w}_{i}$. The plane equation of i-th wall is $a_{i}x+b_{i}y+c_{i}z+d=0$ and $\mathds{1}(\prod_{\mathbf{w}_{i}}(x_{jp},y_{jp},z_{jp})\,{\rm in}\,\mathbf{w}_{i})$ indicates whether the projection of bounding box vertices $(x_{jp},y_{jp},z_{jp})$ of j-th object is inside $\mathbf{w}_{i}$. We skip some objects such as windows and doors since they can intersect with walls. We adopt the 3D IoU loss $\mathcal{L}_{\rm o-o}$ in DiffuScene as follows,
The denoising network $\mathbf{\epsilon_{\theta}}$ takes the scene code $\mathbf{x_{t}}$, text prompt $y$, and timestep $t$ as input, and denoises them iteratively to get a clean scene code $\mathbf{\hat{x}}_{0}$. Please refer to the supplementary file for details of our scene code denoising network.

### v4

@@ Abstract
Text-driven 3D indoor scene generation is useful for gaming, film industry, and AR/VR applications. However, existing methods cannot faithfully capture the scene layout based on text descriptions, nor do they allow flexible editing of individual objects in the room. To address these problems, we present Ctrl-Room, which can generate convincing 3D rooms with designer-style layouts and high-fidelity textures from just a text prompt. Our key insight is to separate the modeling of layouts and appearance. Our proposed method consists of two stages: a Layout Generation Stage and an Appearance Generation Stage. The Layout Generation Stage trains a text-conditional diffusion model to learn the layout distribution with our holistic scene code parameterization. Next, the Appearance Generation Stage employs a fine-tuned ControlNet to produce a vivid panoramic image of the room guided by the 3D scene layout, then further upgrades to a panoramic NeRF model. Benefiting from the scene code parameterization, we can easily edit the generated room model through our mask-guided editing module, without expensive edit-specific training. Extensive experiments on the Structured3D dataset demonstrate that our method outperforms existing methods in producing more reasonable, view-consistent, and editable 3D rooms from text prompts.
@@ 1 Introduction
High-quality 3D indoor scenes play a crucial role across a wide array of applications, ranging from interior design and video games to simulators for embodied AI. Traditionally, indoor scenes are crafted manually by professional artists, which is both time-consuming and costly. Recent advancements in generative models [22, 5, 18, 29] have attempted to simplify the creation of 3D models from textual descriptions, However, extending this capability to text-driven 3D indoor scene generation presents unique challenges as they exhibit strong semantic layout constraints, such as neighboring walls are perpendicular and the TV set often faces a sofa, that are more complicated than objects.
@@
Existing text-driven 3D indoor scene generation approaches, such as Text2-Room [12] and Text2NeRF [46], are designed with an incremental framework. They create 3D indoor scenes by incrementally generating different viewpoints frame-by-frame and reconstructing the 3D mesh of the room from these sub-view images. However, these approaches often fail to model the global layout of the room, resulting in unconvincing results. As shown in the first row of  Fig. 1 (a), the result of Tex2Room exhibits repetitive objects, e.g. several cabinets in a living room, and does not follow the furniture layout patterns. We refer to this problem as the ‘Penrose Triangle problem’, where a generated scene has plausible 3D structures everywhere locally but lacks global consistency. Furthermore, prior approaches do not offer user-friendly interaction, as the resulting 3D geometry and textures are not editable. Other method [16, 17, 30, 36] represent the scene as a panorama image and generate it from a text prompt. However, these works cannot guarantee reasonable scene layouts. As shown on the middle row of Fig. 1 (a), a bedroom generated by MVDiffusion [36] contains multiple beds, which violates room layout priors.
@@
To address these shortcomings, we propose a novel two-stage method to generate a high-fidelity and editable 3D room. The key insight is to separate the generation of 3D geometric layouts from that of visual appearance, which allows us to better capture the room layout and achieve vivid textures at the same time. In the first stage, from text input, our method creates plausible scene layouts with various furniture types and positions. Unlike previous scene synthesis methods [34, 21] that only focus on the furniture arrangement, our approach further considers walls with doors and windows, which play an essential role in the layout. To achieve this goal, we parameterize the room by a holistic scene code, which represents a room as a set of objects. Each object is represented by a vector capturing its position, size, semantic class, and orientation. Based on our compact parameterization, we design a diffusion model to learn the 3D room layout distribution from the Structured3D dataset [48].
@@
Our method then generates the room appearance with the guidance of the 3D room layout. We first generate a panorama using a text-to-image latent diffusion model, then iteratively upgrade the generated images to a NeRF model and generate additional novelty view panorama images. During the panorama generation, unlike previous text-to-panorama works [36, 6], our method explicitly enforces scene layout constraints and guarantees plausible 3D room structures and furniture arrangement. To achieve this goal, we convert the 3D layout synthesized in the first stage into a semantic segmentation map and feed it to a fine-tuned ControlNet [47] model to create the panorama image. We also use this layout information to estimate scene depth and inpaint missing regions at novel viewpoints.
@@
Benefiting from the separation of layout and appearance, our method enables flexible editing on the generated 3D room. The user can replace or modify the size and position of furniture items, e.g. replacing the TV and TV stand as in Fig. 1 (b). Our method can update the room according to the edited room layout through our mask-guided editing module without expensive edit-specific training. The updated room appearance maintains consistency with the original version while satisfying the user’s edits.
@@
  • To address the Penrose Triangle Problem, we design a two-stage method for 3D room generation from pure text input, which separates the geometric layout generation and appearance generation. In this way, our method can better capture the scene layout constraints in real-world data and produce a vivid appearance simultaneously.
  • Within the separated layout and appearance generation, we introduce novel techniques, including holistic scene code parametrization, layout-guided panorama generation, layout-guided panoramic NeRF, and a mask-guided editing module to achieve high-quality and flexible 3D room generation.
  • Qualitative and quantitative experiments confirm that our method excels in producing more realistic and editable 3D rooms compared to existing approaches.
@@ 2 Related Work
Figure 2: Framework overview. In Layout Generation Stage, we synthesize a scene code from the text input and convert it to a 3D bounding box representation to facilitate editing. In Appearance Generation Stage, we project the bounding boxes into a semantic segmentation map to guide the panorama synthesis. The panorama is then reconstructed into a panoramic NeRF (PeRF)[38]model with layout guidance.
@@ 2.1 Text-based 3D Object Generation
Early methods employ 3D datasets to train generative models. Text2Shape [4] learns a feature representation from paired text and 3D data and uses GAN to generate 3D shapes from the text. Point-E [19] and Shap-E [13] enlarge the scope of the training dataset and employ a latent diffusion model [24] for object generation. However, 3D datasets are scarce, which makes these methods difficult to scale. More recent methods [19, 22, 18, 39, 5, 41] exploit the powerful 2D text-to-image diffusion models [24, 26] for 3D model generation. Typically, these methods generate one or multiple 2D images in an incremental fashion and optimize the 3D model accordingly. DreamFusion [22] introduces a loss based on probability density distillation and optimizes a randomly initialized 3D model through gradient descent. Magic3D [18] uses a coarse model to represent 3D content and accelerates it using a sparse 3D hash grid structure. To alleviate over-saturation and low-diversity problems, ProlificDreamer [41] models and optimizes the 3D parameters through variational score distillation. However, these methods are limited to 3D object generation and cannot be directly extended to 3D scene generation which has additional layout constraints.
@@ 2.2 Text-based 3D Room Generation
Layout generation has been greatly boosted by transformer-based methods. LayoutTransformer [10] employs self-attention to capture relationships between elements to accomplish layout completion. ATISS [21] proposes an autoregressive transformer to generate proper indoor scenes with only the room type and floor plan as the input. DiffuScene [34] and InstructScene [15] model a union of furniture as a fully connected scene graph and propose a diffusion model to sample physically plausible scenes. While these methods generate reasonable furniture layouts, they do not consider the walls, doors, and windows which are crucial in the furniture arrangement. Thus they do not always generate realistic indoor environments.
@@
Another line of work [16, 17, 30] represent an indoor scene by a panorama image without modeling 3D shapes. These methods enjoy the benefits of abundant training data and produce vivid results. COCO-GAN [16] produces a set of patches and assemble them into a panoramic image. InfinityGAN [17] uses the information of two patches to generate the parts between them, to finally obtain a panoramic image.  [30] proposes a 360-aware layout generator to produce furniture arrangements and uses this layout to synthesize a panoramic image based on the input scene background. MVDiffusion [36] simultaneously generates multi-view perspective images and proposes a correspondence-aware attention block to maintain multi-view consistency, and then transfers these images to a panorama. These methods might suffer from incorrect room layout since they do not enforce layout constraints. Furthermore, the results of these methods cannot be easily edited, e.g. resizing or moving furniture around, because they do not maintain an object-level representation.
@@
GAUDI [2] generates immersive 3D indoor scenes rendered from a moving camera. It disentangles the 3D representation and camera poses to ensure the consistency of the scene during camera movement. CC3D [1] proposes a 3D-aware GAN for multi-object scenes conditioned on a single semantic layout image and is trained using posed multi-view RGB images. Another related line of work [32, 43, 28] deals with retexturizing a given 3D scene. They employ 2D diffusion models to stylize and further improve the given geometry. Text2Room [12] incrementally synthesizes nearby images with a 2D diffusion model and recovers its depth maps to assemble into a 3D room mesh. Unfortunately, it cannot handle the geometric and textural consistency among the images, resulting in the ‘Penrose Triangle problem’. In our method, we take both geometry and appearance into consideration and create a more geometrically plausible 3D room. A concurrent work [28] also guides the 3D room mesh generation by leveraging the user-input scene layouts. In contrast, our method is capable of synthesizing professional designer-style layouts solely from text prompts.
@@ 3 Method
In order to achieve text-based 3D indoor scene generation, we propose Ctrl-Room. We first generate the room layout from an input text and then generate the room appearance represented by panoramic images according to the layout, followed by layout-guided panoramic NeRF [38] to generate the final 3D room. This mechanism solves the Penrose Triangle Problem to generate physically plausible 3D rooms, while also enabling users to edit the scene layout interactively. The overall framework of our method is depicted in  Fig. 2, which consists of two stages: the Layout Generation Stage and the Appearance Generation Stage. In the Layout Generation Stage, we parameterize the indoor scene with a holistic scene code and design a diffusion model to learn its distribution. Once the holistic scene code is generated from text, we recover the room as a set of orientated bounding boxes of walls and objects. Note that users can edit these bounding boxes by adjusting their semantic types, positions, or scales, enabling the customization of 3D room generations. In the Appearance Generation Stage, we obtain an RGB panorama through a conditioned image diffusion model to represent the room texture. Specifically, we project the generated layout bounding boxes into a semantic segmentation layout. We then fine-tune a pre-trained ControlNet [47] model to generate an RGB panorama from the input semantic layout. To ensure loop consistency, we propose a loop-consistent sampling during the inference process. Finally, we integrate the layout and the panorama, then generate a full 3D room through the layout-guided panoramic NeRF module [38]. This module progressively inpaints panoramas at new viewpoints using the fine-tuned ControlNet. To extract meshes from reconstructed NeRF, we render depth maps of the new views and utilize truncated signed distance fusion (TSDF) to obtain the final mesh.
@@ 3.1 Layout Generation Stage
Scene Code Definition. Different from previous methods  [21, 34], we consider not only furniture but also walls, doors, and windows to define the room layout. We employ a unified encoding of various objects. Specifically, given a 3D scene $\mathcal{S}$ with $m$ walls and $n$ furniture items, we represent the scene layout as a holistic scene code $\mathbf{x_{0}}=\{\mathbf{o_{i}}\}^{N}_{i=1}$, where $N=m+n$. We encode each object $o_{j}$ as a node with attributes including center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. The concatenation of these attributes characterizes each node as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},r_{i}]$. As can be seen in Fig. 3 (a), we represent a scene layout as a tensor $\mathbf{x_{0}}\in\mathbb{R}^{N\times D}$, where $D$ is the attribute dimension of a node. In all the data, we choose the normal direction of the largest wall as the ‘main direction’. For other objects, we take the angles between their front directions and the main direction as their rotations. We use the one-hot encoding to represent their semantic types.
@@
where $\epsilon_{\theta}$ is the noise estimator which aims to find the noise $\mathbf{\epsilon}$ added into the input $x_{0}$. Here, $y$ is the text embedding of the input text prompts. The hyperparamter $w_{t}$ is set to $\mathbf{\bar{\alpha}_{t}}*0.1$. $\mathcal{L}_{\rm w-o}$ is the physical violation loss between walls and objects. We adopt the 3D IoU loss $\mathcal{L}_{\rm o-o}$ in DiffuScene to avoid intersection between furniture.
The denoising network $\mathbf{\epsilon_{\theta}}$ takes the scene code $\mathbf{x_{t}}$, text prompt $y$, and timestep $t$ as input, and denoises them iteratively to get a clean scene code $\mathbf{\hat{x}}_{0}$. Please refer to appendix Sec.1 for the details of our $\mathcal{L}_{\rm w-o}$ and denoising network.